OpenStudy (anonymous):

Anyone know what this would be..? -5c/8 = 20

OpenStudy (solomonzelman):

sure:)

OpenStudy (solomonzelman):

\(\large\color{slate}{ \displaystyle \frac{-5c}{8} = 20 }\) is same as \(\large\color{slate}{ \displaystyle \frac{-5}{8}c = 20 }\)

OpenStudy (solomonzelman):

multiply both sides times \(\large\color{slate}{ \displaystyle \frac{8}{-5} }\)

OpenStudy (anonymous):

Oh K thxs so much...! :)

OpenStudy (solomonzelman):

yw, glad that even the smalest hint already helped you out.

OpenStudy (anonymous):

Lol...! :)

OpenStudy (solomonzelman):

:)